LgswUnitSx and LgswUnitDx HW servicing

8:00: we unpack the PCD and the amplifier from 3L boxes, we bring it to 5. We also spend 1:30h in searching for the WfsCameraPump1 to evacuate the PnCCD.

Jose applies a patch to LGSW housekeeping: now THORs on SX take about 15s to reconnect after booting. He also fixes Zabbix (several sensors where missing).

10:00: start pumping PnCCD. In the meanwhile we install the PCD and ampli on DX-LGSW (it has a different IP, Julian sent us an email about HowToChangePockelsCellAmplifierIp).

We keep the pump on for about 2h, at the end the gauge reads 0.35 mbar. After 5min it reads 0.4mbar...

AOS v15.5 check w/ AARB

While pumping we check for the new AOS version.

Note: to login as tcs@tcs1 we must first login as telescope on an OBS machine.

We login as argos on FLAO on both DX and SX. The AARBs are on. On DX we have to revert to the default config file following instructions from Luca. After this all sides seem to be fine.

  • On DX the AOS GUI says the LBTIWFS is connected

IMG_20160511_122808560_HDR.jpg

New LAT focus check

12:00: we move to horizon, Gustavo must be trained to install and remove the AGW cubes.

While at horizon we also check for the LAT power supply indeed the basdard cannot connect to it. Power and ethernet to the MOXA seem fine. After a couple of checks Jose try to connect to it through the LATClient. He sends a CamPowerOn command and the LAT basdard turns green.
  • We must dig out why the basdard cannot connect to LAT at fresh start.

IMG_20160511_133326277.jpg

In the meanwhile Jose also fixed a problem with the SVN repository: Florian changed the SVN server, the repository had a new id for several hours, then he hacked it to get back the old id...

Sebastian and WG check for new LAT focus stage...

14:00: LBC filling, it takes about 1h.

Test of PockelsCellsTrasmission

14:45: PCT on SX-LGSW:

20160511_213750_pcTransmission.jpg

CL test of LGSW + PYR BIN2

15:00: back to Zenith, we deploy the cal swing arm. Luca has restored the old AOS version.

We have problems in finding light because the tertiary selector mirror didn't home properly. Once fixed we get back light with hexapod in the March position:

SX HEXA 5.20 0.72 -0.75 136 -39
The SX system was started from arbitrator, FLAO operated from argos account. TS service started properly.

We first close the loop with LGSW only, we center the FLAO stages on the on-axis source position.

SX BOARD -73.8 -52.6 45.0

  • The TS seems to not work properly, ERROR addind temporary modes is displayed in the TS log. Restarting the process fixes this.

Then we close the loop with Pyr + LGSW. On axis source equivalent to mag 12 star. PyrWfs set to Bin2 990 Hz. Truth sensing is integrating 990 frames (1s integration).

We change the on-axis source to mag 14. PyrWfs set at Bin4 300Hz. TS integrates 300 frames (1s integration).

We switch to QC, keeping mag 14. PyrWfs is set to Bin4 100Hz. TS integrates 500 frames equivalent to 5s integration.

-- MarcoBonaglia - 11 May 2016
Topic attachments
I Attachment Action Size Date Who Comment
20160511_213750_pcTransmission.jpgjpg 20160511_213750_pcTransmission.jpg manage 54 K 11 May 2016 - 21:56 UnknownUser Pockel Cell Transmission Test
IMG_20160511_122808560_HDR.jpgjpg IMG_20160511_122808560_HDR.jpg manage 1 MB 11 May 2016 - 22:02 UnknownUser  
IMG_20160511_133326277.jpgjpg IMG_20160511_133326277.jpg manage 3 MB 11 May 2016 - 22:08 UnknownUser  
Topic revision: r7 - 23 May 2016, LorenzoBusoni
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback